OpenStudy (anonymous):

What is the energy of a single photon of visible light with a wavelength of 500.0 nm? 1 nm = 10^-9 m (Points : 5) 2.02 × 10^18 J 2.84 × 10^14 J 3.97 × 10^-19 J 4.95 × 10^-19 J

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

here we have to use this formula: \[E = h\nu = \frac{{hc}}{\lambda }\] where h is the Planck's constant, namely h=6.62*10^(-34) Joules*sec

OpenStudy (michele_laino):

\lambda is the wave length, and c= 3*10^8 m/sec is the light speed